OpenStudy (anonymous):

Find the points on r = a (1 + cosθ), a > 0 , where the tangent line is horizontal.

OpenStudy (dumbcow):

The tangent line is horizontal when max or min occur for "y" normally set dy/dx = 0 since its in polar we can use dy/dtheta = 0 which will find the angles where max/min occur \[y = r \sin \theta = a(1+\cos \theta) \sin \theta\] \[\frac{dy}{d \theta} = a (-\sin^2 \theta +\cos \theta + \cos^2 \theta) = 0\] \[\rightarrow 2 \cos^2 \theta + \cos \theta -1 = 0\] \[\rightarrow (2 \cos \theta -1) (\cos \theta +1) = 0\] \[\cos \theta = \frac{1}{2} \rightarrow \theta = \pm \frac{\pi}{3}\] \[\cos \theta = -1 \rightarrow \theta = \pi\]